OpenStudy (anonymous):

Show that sin(2x) =2t/(1+t^2) when t = tan x

OpenStudy (anonymous):

Sin(2x)=2Sin(x)Cos(x) and 1+(tanx)^2= (Secx)^2 ..Also Secx=1/Cosx ...This problem becomes quite simple after these substitutions are made....therefore 2tanx/1+(tanx)*2 ends up being 2SinxCosx which is Sin(2x)....Hope this helps...